Simplify. 225+25 (2) -
Answer #1

Solution:-

Given that,

2\sqrt{3}+2\sqrt{3}(\sqrt{2})-\sqrt{3}

Let's simplify the above expression

2\sqrt{3}+2\sqrt{3}(\sqrt{2})-\sqrt{3}

Removing parentheses:

  • (a) = a

Therefore,

2\sqrt{3}+2\sqrt{3}(\sqrt{2})-\sqrt{3}

2\sqrt{3}+2\sqrt{3}\sqrt{2}-\sqrt{3}

Adding similar elements:

2\sqrt{3}-\sqrt{3}+2\sqrt{3}\sqrt{2}

\sqrt{3}+2\sqrt{3}\sqrt{2}

Applying radical rule:

  • \sqrt{a}\sqrt{b} = \sqrt{a\times b}

Hence,

\sqrt{3}+2\sqrt{3}\sqrt{2}

\sqrt{3}+2\sqrt{3\times 2}

\boldsymbol{\sqrt{3}+2\sqrt{6}}

Hence,

2\sqrt{3}+2\sqrt{3}(\sqrt{2})-\sqrt{3} = \boldsymbol{\sqrt{3}+2\sqrt{6}}\; \; \; \; \; (Answer)